# FABG : End-to-end Imitation Learning for Embodied Affective Human-Robot Interaction

### v1

@@ Abstract
This paper proposes FABG (Facial Affective Behavior Generation), an end-to-end imitation learning system for human-robot interaction, designed to generate natural and fluid facial affective behaviors. In interaction, effectively obtaining high-quality demonstrations remains a challenge. In this work, we develop an immersive virtual reality (VR) demonstration system that allows operators to perceive stereoscopic environments. This system ensures ”the operator’s visual perception matches the robot’s sensory input” and ”the operator’s actions directly determine the robot’s behaviors” — as if the operator replaces the robot in human interaction engagements. We propose a prediction-driven latency compensation strategy to reduce robotic reaction delays and enhance interaction fluency. FABG naturally acquires human interactive behaviors and subconscious motions driven by intuition, eliminating manual behavior scripting. We deploy FABG on a real-world 25-degree-of-freedom (DoF) humanoid robot, validating its effectiveness through four fundamental interaction tasks: expression response, dynamic gaze, foveated attention, and gesture recognition, supported by data collection and policy training. Project website: https://cybergenies.github.io
@@ I INTRODUCTION
Humanoid robots have been increasingly deployed in human-centric scenarios, including critical domains such as educational assistance, behavioral guidance, and emotional companionship [6, 1]. Research demonstrates that nonverbal behaviors — particularly postural adjustments, gestural dynamics, and facial micro-expressions — play an essential role in conveying internal states during human interactions [14, 27]. Within this context, imitation learning (IL), as a prominent reinforcement learning paradigm, demonstrates significant advantages by enabling robots to acquire skills through expert behavior emulation [36]. Existing studies have validated IL’s effectiveness in robotics, where robots achieve task execution ranging from simple manipulations to multi-step complex operations via human demonstration observation [11, 24, 28, 34]. These capabilities establish IL as a promising approach for driving autonomous affective interaction in humanoid robots. However, this presents two core challenges: how to efficiently acquire high-quality facial expression and motion demonstration data, and how to design imitation learning strategies suitable for facial expression interaction scenarios.
@@
Current mainstream demonstration data acquisition methods rely on robotic teleoperation systems. Although existing studies have adopted interactive tools such as VR/AR controllers [3, 15, 9], 3D spatial mice [7, 37], smartphones [17], and haptic devices [29, 20], these solutions universally exhibit practical limitations: high-precision equipment incurs substantial costs, control interfaces demonstrate notable latency, and user interaction logic lacks intuitiveness. Regarding teleoperation perception, some systems directly utilize the operator’s eyes to observe the robot’s task space, resulting in the operator’s view being frequently occluded by the robot’s body, thereby hindering real-time evaluation of action execution outcomes.
@@
As a crucial paradigm for robotic skill acquisition, imitation learning methods have achieved remarkable progress in operational task learning [13, 2, 33]. The recently proposed Action Chunking with Transformers (ACT) [35] enhances complex manipulation task performance through a task decomposition mechanism that segments long-horizon tasks into manageable temporal chunk sequences, significantly improving learning efficiency and execution coherence. However, ACT exhibits two critical limitations: (i) the original model relies solely on RGB inputs without leveraging depth information, and (ii) its Temporal Ensemble suffers from historical error accumulation in real-time interaction scenarios while disabling Temporal Ensemble leads to action discretization and execution stuttering.
@@
To address these challenges, we propose FABG — an end-to-end system for embodied affective interaction — as shown in Figure 1, with three core technical breakthroughs: First, an immersive VR teleoperation system. By designing a dual-modality data acquisition architecture that integrates ZED stereo cameras into VR devices, we synchronously capture the robot’s first-person visual streams with the operator’s facial landmarks/head pose data, enabling efficient acquisition of high-quality human demonstrations. Second, depth-enhanced observation representation. We extend ACT’s RGB input to RGB-D multimodal input, where the depth channel enhances spatial comprehension of human body language through dynamic distance awareness. Third, a prediction-driven latency compensation strategy that achieves triple optimization via prospective prediction: mitigating historical error propagation, eliminating mechanical jitter from discrete actions, and compensating multi-source system delays.
Our principal contributions are:
@@ III-A Immersive VR Demonstration System
Employing appropriate system for effective demonstration collection is critical to obtaining high-quality results. In this work, we present a virtual reality-based immersive system for collecting human user demonstrations. As shown in Fig. 1, human demonstrators utilize PICO 4 Pro VR headsets and ZED RGB-D cameras to perform interactive behaviors instead of physical robots. Compared to using two independent biomimetic cameras in robotic eyes, ZED RGB-D cameras provide stable and precise depth information. The VR system records the operator’s 58 ARKit facial expression coefficients and head RPY (Roll, Pitch, Yaw)
rotational movements in real-time, while the RGB-D camera synchronously captures 480×640 resolution first-person-view imagery. To equip operators with active first-person perception capabilities and align human-robot viewpoints, we developed a 3D environmental perception module in Unity, implementing stereoscopic passthrough functionality that enables operators to observe 3D environments consistent with the real world. As depicted in Figure 2, we calibrated the position and scaling of ZED camera feeds within the VR interface to achieve seamless stitching between 3D passthrough views and ZED camera fields of view. Spatial markers were then applied to the ZED video streams before disabling their display, thereby guiding operators into the correct visual zones.
@@ III-C Prediction-Driven Latency Compensation
While ACT achieves policy optimization through Action Chunking and Temporal Ensemble mechanisms, it exhibits significant limitations in natural human-robot interaction. In the implementation of Action Chunking, the agent generates a k-frame action sequence at timestep for execution until . This discrete decision-making causes trajectory discontinuities between boundary frames ( and ) of adjacent chunks due to unoptimized motion continuity across blocks, manifesting as periodic motion stuttering in robots. The root cause lies in the chunked generation strategy disrupting the temporal consistency of action flows. Although Temporal Ensemble mitigates stuttering by querying policies at every timestep, it introduces dual challenges: its reliance on accumulated historical predictions induces delayed responses to immediate commands while amplifying error accumulation from environmental noise. Compounded by inherent system delays (perception-actuation latency, computational latency, communication latency, etc.), these factors exacerbate response lag and severely degrade interaction quality.
To address these issues, we propose Prediction-Driven Latency Compensation (PDLC) as shown in Figure 3. A three-phase coordinated mechanism enhances performance: First, we integrate a full-temporal action prediction framework into conventional chunking, where multi-step action sequences are generated at every timestep to eliminate temporal window abruptness in fixed chunking through high-frequency prediction. Second, we design a dynamic truncation mechanism based on predefined temporal offsets, determining optimal action truncation offsets via systematic delay quantification to achieve end-to-end latency compensation. Finally, an implicit continuity preservation mechanism coordinates trajectory continuity across prediction windows through synergistic dynamic action truncation and policy iteration.
@@ IV-A Hardware Platform
We validate human-robot interaction behaviors using a self-designed bio-inspired robotic head platform that simulates human-like facial expressions and head motions through a rigid-flexible hybrid structure and multi-DoF actuation system. The core architecture comprises a 3D-modeled rigid framework integrated with elastic bio-inspired skin. By converting expression parameters into 25-channel PWM signals to drive actuators, we achieve independent/coordinated control of key facial regions including eyelids, eyebrows, cheeks, and mouth. This platform establishes a controllable experimental interface for affective computing and behavioral mirroring research via engineering mapping of biomechanical characteristics.
@@ IV-B Experimental Setup
Affective Interaction Task: Requires real-time facial expression recognition and congruent response generation. The robot must reciprocate smiles with smiles and surprise expressions with corresponding reactions. When testers exhibit subtle expression intensity variations, the robot synchronously adjusts its responses while maintaining mirroring during static expressions. Testers perform 10 smiling and 10 surprised expressions.
@@
Foveated Attention Task: Tests dynamic visual pursuit capabilities. Testers perform finger approach-retraction motions, during which the robot adaptively adjusts inter-pupillary distance (IPD) to maintain gaze focus during proximity and resets IPD upon retraction. This task requires precise distance sensitivity for continuous finger tracking via dynamic IPD adaptation. 10 trials are conducted.
Gesture Recognition Task: Assesses dynamic command parsing. Testers maintain hand positions while randomly cycling through five gestures (up/down/left/right/fist). The robot must accurately recognize gestures to reorient toward specified directions or fixate on fists. 20 trials (4 per gesture) are executed.
@@ IV-C Experimental Results
The input modalities cover both RGB and RGB-D configurations, while the execution strategies encompass three approaches: without temporal ensemble (w/o TE), temporal ensemble (TE), and PDLC, resulting in six comparative experimental groups. All models are built upon the DinoV2 [19] visual backbone and trained under unified computational conditions (NVIDIA RTX 4070Ti GPU) with identical hyperparameters: initial learning rate , batch size 20, and 20,000 optimization iterations. To eliminate computational bias, all experiments execute inference at 30Hz. Validation is conducted on the facial robot platform described in Section IV-A, ensuring environmental consistency.
@@
As per Table I’s five-trial averaged results, our proposed PDLC strategy with RGB-D input demonstrates superior performance. In affective interaction tasks, models without inference strategies fail due to state instability, while RGB-D+PDLC achieves 6.94s (smile) and 7.55s (surprise) response times — showing 84.8 and 85.9 improvements over RGB-D+TE. Depth sensing enhances micro-expression capture, reducing surprise response time by 37.2 compared to RGB+PDLC. For dynamic tracking, PDLC’s motion trajectory prediction improves continuity, with RGB-D+PDLC completing hand tracking in 6.77s (37.0 faster than RGB+PDLC), confirming depth-enhanced 3D spatial localization. In foveated attention tasks, RGB-D+PDLC adjusts inter-pupillary distance in 11.44s (35.4 reduction vs. RGB+PDLC), benefiting from depth-based precise distance perception. Gesture recognition tasks show RGB-D+PDLC resolving commands in 14.23s (85.6 faster than TE), attributed to PDLC’s temporal noise suppression. TE strategies increase average task durations by 21-25 due to historical interference, whereas PDLC reduces average time by 82.6 via dynamic compensation. Multi-group comparisons confirm that RGB-D and PDLC synergistically excel in spatial-sensitive tasks, providing robust solutions for real-time HRI systems.
@@ IV-D Ablation experiment
To validate the effectiveness of PDLC in achieving low-latency response and high motion accuracy, this study designed a facial expression imitation experiment. Participants wearing motion capture devices executed standardized action sequences including sustained mouth-opening and rapid open-close motions, while their facial movements and the robot’s response trajectories were synchronously recorded. Within the FABG system, a model with RGB-D input modality was employed to comparatively evaluate three configurations: the strategy without temporal reasoning (w/o TE), the temporal ensemble strategy (TE), and the PDLC strategy.

### v2

@@ Abstract
This paper proposes FABG (Facial Affective Behavior Generation), an end-to-end imitation learning system for human-robot interaction, designed to generate natural and fluid facial affective behaviors. In interaction, effectively obtaining high-quality demonstrations remains a challenge. In this work, we develop an immersive virtual reality (VR) demonstration system that allows operators to perceive stereoscopic environments. This system ensures that ”the operator’s visual perception matches the robot’s sensory input” and ”the operator’s actions directly determine the robot’s behaviors” - as if the operator replaces the robot in human interaction engagements. We propose a prediction-driven latency compensation strategy to reduce robotic reaction delays and enhance interaction fluency. FABG naturally acquires human interactive behaviors and subconscious motions driven by intuition, eliminating manual behavior scripting. We deploy FABG on a real-world 25 degree-of-freedom (DoF) humanoid robot, validating its effectiveness through four fundamental interaction tasks: affective interaction, dynamic tracking, foveated attention, and gesture recognition, supported by data collection and policy training.
@@ I INTRODUCTION
Humanoid robots have been increasingly deployed in human-centric scenarios, including critical domains such as educational assistance, behavioral guidance, and emotional companionship [6, 1]. Research shows that nonverbal behaviors - particularly postural adjustments, gestural dynamics, and facial microexpressions - play an essential role in the transmission of internal states during human interactions [14, 27]. Within this context, imitation learning (IL), as a prominent reinforcement learning paradigm, demonstrates significant advantages by enabling robots to acquire skills through expert behavior emulation [36]. Existing studies have validated the effectiveness of IL in robotics, where robots achieve task execution ranging from simple manipulations to multi-step complex operations via human demonstration observation [11, 24, 28, 34]. These capabilities establish IL as a promising approach for driving autonomous affective interactions in humanoid robots. However, this presents two core challenges: how to efficiently acquire high-quality facial expression and motion demonstration data, and how to design imitation learning strategies suitable for facial expression interaction scenarios.
@@
Current mainstream demonstration data acquisition methods are based on robotic teleoperation systems. Although existing studies have adopted interactive tools such as VR/AR controllers [3, 15, 9], 3D spatial mice [7, 37], smartphones [17], and haptic devices [29, 20], these solutions universally exhibit practical limitations: high-precision equipment incurs substantial costs, control interfaces demonstrate notable latency, and user interaction logic lacks intuitiveness. Regarding teleoperation perception, some systems directly utilize the operator’s eyes to observe the robot’s task space, resulting in the operator’s view being frequently occluded by the robot’s body, thereby hindering real-time evaluation of action execution outcomes.
@@
As a crucial paradigm for robotic skill acquisition, imitation learning methods have achieved remarkable progress in operational task learning [13, 2, 33]. The recently proposed Action Chunking with Transformers (ACT) [35] enhances the performance of complex manipulation tasks through a task decomposition mechanism that segments long-horizon tasks into manageable temporal chunk sequences, significantly improving learning efficiency and execution coherence. However, ACT exhibits two critical limitations: (i) the original model relies solely on RGB inputs without leveraging depth information, and (ii) its Temporal Ensemble suffers from historical error accumulation in real-time interaction scenarios while disabling Temporal Ensemble leads to action discretization and execution stuttering.
@@
To address these challenges, we propose FABG - an end-to-end system for embodied affective interaction - as shown in Figure 1, with three key technical breakthroughs: First, an immersive VR teleoperation system. By designing a dual-modality data acquisition architecture that integrates ZED stereo cameras into VR devices, we synchronously capture the robot’s first-person visual streams with the operator’s facial landmarks/head pose data, enabling efficient acquisition of high-quality human demonstrations. Second, a depth-enhanced observation representation. We extend ACT’s RGB input to RGB-D multimodal input, where the depth channel enhances spatial comprehension of human body language through dynamic distance awareness. Third, a prediction-driven latency compensation strategy that achieves triple optimization via prospective prediction: mitigating historical error propagation, eliminating mechanical jitter from discrete actions, and compensating multi-source system delays.
Our principal contributions are as follows:
@@ III-A Immersive VR Demonstration System
Employing appropriate system for effective demonstration collection is critical to obtaining high-quality results. In this work, we present a virtual reality-based immersive system for collecting human user demonstrations. As shown in Fig. 1, human demonstrators utilize PICO 4 Pro VR headsets and ZED RGB-D cameras to perform interactive behaviors instead of physical robots. Compared to using two independent biomimetic cameras in robotic eyes, ZED RGB-D cameras provide stable and precise depth information. The VR system records the operator’s 58 ARKit facial expression coefficients and head RPY (Roll, Pitch, Yaw) rotational movements in real-time, while the RGB-D camera synchronously captures 480×640 resolution first-person-view imagery. To equip operators with active first-person perception capabilities and align human-robot viewpoints, we developed a 3D environmental perception module in Unity, implementing stereoscopic passthrough functionality that enables operators to observe 3D environments consistent with the real world. As depicted in Figure 2, we calibrated the position and scaling of ZED camera feeds within the VR interface to achieve seamless stitching between 3D passthrough views and ZED camera fields of view. Spatial markers were then applied to the ZED video streams before disabling their display, thereby guiding operators into the correct visual zones.
@@ III-C Prediction-Driven Latency Compensation
ACT generates precise and smooth motions through action chunking and temporal ensemble. However, this approach exhibits significant limitations in real-time interaction. When using action chunking alone, the robot generates a sequence of frames at time and executes them until time . This discrete decision-making mechanism results in trajectory discontinuities at the junctions of adjacent action sequences, leading to periodic motion stuttering. The temporal ensemble avoids this issue by querying the policy at each time step. However, because it accumulates historical predictions, it inherently introduces response delays to immediate commands, potentially leading to error accumulation and amplification in dynamic environments due to noise interference. Moreover, multiple inherent system delays, such as the operator’s perception-action latency during data acquisition, computational delays in inference, and communication delays, further exacerbate response lag, ultimately impairing the real-time interaction experience between humans and robots.
To address these issues, this study proposes Prediction-Driven Latency Compensation (PDLC). Our approach queries the policy at each time step to prevent periodic motion stuttering. Additionally, we design a dynamic truncation mechanism based on predefined time offsets. By quantitatively analyzing system delays, we determine the optimal action offset to achieve effective end-to-end latency compensation. Specifically, we fix the length of an action sequence to , with representing the optimal action offset. At each time step, the robot receives an observation, generates an action sequence of length , and executes the -th action from the sequence, as illustrated in Figure 3. This process incurs no additional training cost or inference time. Experimental results demonstrate that PDLC is crucial for improving real-time human-robot interaction, as it significantly reduces response latency and generates precise and smooth interactive motions.
@@ IV-A Hardware Platform
We validate human-robot interaction behaviors using a self-designed bioinspired robotic head platform that simulates human-like facial expressions and head motions through a rigid-flexible hybrid structure and a multi-DoF actuation system. The core architecture comprises a 3D-modeled rigid framework integrated with elastic bioinspired skin. By converting expression parameters into 25-channel PWM signals to drive actuators, we achieve independent/coordinated control of key facial regions including eyelids, eyebrows, cheeks, and mouth. This platform establishes a controllable experimental interface for affective computing and behavioral mirroring research through engineering mapping of biomechanical characteristics.
@@ IV-B Experimental Setup
Affective Interaction Task: Requires real-time facial expression recognition and congruent response generation. The robot must reciprocate smiles with smiles and surprise expressions with corresponding reactions. When testers exhibit subtle variations in the intensity of expression, the robot synchronously adjusts its responses while maintaining mirroring during static expressions. The testers present 10 smiling expressions and 10 surprised expressions.
@@
Foveated Attention Task: Tests dynamic visual pursuit capabilities. The testers perform finger approach-retraction motions, during which the robot adaptively adjusts the interpupillary distance (IPD) to maintain gaze focus during proximity and resets IPD upon retraction. This task requires precise distance sensitivity for continuous finger tracking via dynamic IPD adaptation. Ten trials are conducted.
Gesture Recognition Task: Assesses dynamic command parsing. The testers maintain hand positions while randomly cycling through five gestures (up/down/left/right/fist). The robot must accurately recognize gestures to reorient toward specified directions or fixate on fists. 20 trials (4 per gesture) are executed.
@@ IV-C Experimental Results
The input modalities cover both RGB and RGB-D configurations, while the execution strategies encompass three approaches: without temporal ensemble (w/o TE), temporal ensemble (w/ TE), and PDLC, resulting in six comparative experimental groups. All models are built upon the DinoV2 [19] visual backbone and trained under unified computational conditions (NVIDIA RTX 4070Ti GPU) with identical hyperparameters: initial learning rate , batch size 20, and 20,000 optimization iterations. To eliminate computational bias, all experiments execute inference at 30Hz. Validation is conducted on the facial robot platform described in Section IV-A, ensuring environmental consistency.
@@
According to Table I’s five-trial averaged results, our proposed PDLC strategy with RGB-D input demonstrates superior performance. In affective interaction tasks, models without inference strategies fail due to state instability, while RGB-D+PDLC achieves 6.94s (smile) and 7.55s (surprise) response times, showing 84.8 and 85.9 improvements over RGB-D+TE. Depth sensing enhances micro-expression capture, reducing surprise response time by 37.2 compared to RGB+PDLC. For dynamic tracking, PDLC’s motion trajectory prediction improves continuity, with RGB-D+PDLC completing hand tracking in 6.77s (37.0 faster than RGB+PDLC), confirming depth-enhanced 3D spatial localization. In foveated attention tasks, RGB-D+PDLC adjusts inter-pupillary distance in 11.44s (35.4 reduction vs. RGB+PDLC), benefiting from depth-based precise distance perception. Gesture recognition tasks show RGB-D+PDLC resolving commands in 14.23s (85.6 faster than TE), attributed to PDLC’s temporal noise suppression. TE strategies increase average task durations by 21-25 due to historical interference, whereas PDLC reduces average time by 82.6 via dynamic compensation. Multi-group comparisons confirm that RGB-D and PDLC synergistically excel in spatial-sensitive tasks, providing robust solutions for real-time HRI systems.
@@ IV-D Ablation experiment
To validate the effectiveness of PDLC in achieving low-latency response and high motion accuracy, this study designed a facial expression imitation experiment. Participants wearing motion capture devices executed standardized action sequences including sustained mouth-opening and rapid open-close motions, while their facial movements and the robot’s response trajectories were synchronously recorded. Within the FABG system, a model with RGB-D input modality was employed to comparatively evaluate three configurations: the strategy without temporal reasoning (w/o TE), the temporal ensemble strategy (w/ TE), and the PDLC strategy.